OpenStudy (anonymous):

Divide $ 104 among A,B & c. In the ratio 1/2; 1/3; 1/4

OpenStudy (anonymous):

multiply each ratio by 12, we get 6:4:3 sum of ratios=6+4+3=13 \[ share of first=\frac{ 6 }{13 }\times 104=48 $\] similarly calculate 2nd & 3rd share.